# Word-specific tonal realizations in Mandarin

### v1

@@ 3.2 Representing form: pitch vectors
For the DLM to implement mappings between form and meaning, every form vector input to the model has to have the same number of dimensions as all the others. However, because we took F0 measurements every 15 ms and the tokens in our data vary in duration, our tokens also vary in the number of measurement points. This means that the raw measurements cannot be used to create the form vectors. The raw pitch contours also have the problem that there are gaps due to voicelessness (cf. Appendix A, Figure A4). To overcome these two problems, we used the GAMs described in Section 2.6 to obtain smoothed pitch contours from which we could extract a standard number of measurements.181818 The smoothing for intervocalic voiceless segments is motivated also by the consideration that the control parameters for pitch realization vary smoothly over time in the voiceless intervals. We generated two predicted pitch contours for each token, one using predictions from the word GAM, and the other using predictions from the corresponding omnibus-segment GAM. Each of these predicted contours was then used to generate F0 predictions at 50 equally spaced time points ranging between 0 and 1 for every individual token.
@@
A speaker’s gender and individual characteristics such as vocal tract anatomy, idiolect, and emotional state at the time of speaking, all have strong effects both on their baseline pitch and on pitch range. Similarly, in both our word GAM and in the corresponding omnibus-segment GAM (Section 2.6), the intercepts are largely dependent on the speaker’s gender and individual identity, and differences in amplitude are largely dependent on speech rate, which we take to reflect both the speakers’ idiolect and their emotional state at the time of speaking, amongst other things. On the semantic side, in normal spoken interaction between humans, a speaker’s identity and emotional state not only contribute to the pitch contours they produce, but are also conceptually available to their interlocutors. In contrast, the CEs used as semantic representations in our DLM modeling (Section 3.1) are based entirely on written text and therefore encode much less information about the speaker. To control for this discrepancy, we centered and scaled the predicted F0 values by token; that is, for each token in our data, and for each GAM, we calculated the mean and range of the 50 predicted F0 values, subtracted the mean from each predicted value, and divided the result by the range. In practice, this means that every token contributed equally to the model fit, irrespective of its baseline pitch or amplitude; without scaling, tokens with a greater amplitude would be taken into account more than those with a lower amplitude. A consequence of the way we centered and scaled the pitch vectors is that our DLM production models generate predictions for the geometric shapes of the contours, but not for absolute pitch or amplitude. [^19] Geometric shape can be defined as ‘all the geometrical information that remains when location, scale, and rotational effects are filtered out from an object’ (Kendall, , 1977).
@@ 3.3.1 Method
We used two different methods to map our pitch vectors onto our semantic vectors in a comprehension network. The first method involves a straightforward linear mapping using the Linear Discriminative Learning (LDL) engine of the DLM. This is equivalent to the standard linear mappings used in statistics for multivariate multiple regression (see, e.g., Gahl and Baayen, , 2024; Heitmeier et al., , 2021, 2024, for introductions). The second method (henceforth ResLDL) complements the linear mapping with an additional deep mapping, making it possible to accommodate nonlinear relations while keeping the model relatively interpretable. ResLDL augments an LDL mapping with a nonlinear deep network, which is given the task of capturing any systematicities that are left unexplained in the residuals of the linear network (hence the name ResLDL). Using both these methods, and comparing the results, allowed us to shed light on the complexity of the relationship between our pitch vectors and our semantic vectors. A more detailed introduction to LDL and ResLDL is provided in Appendix C.
@@
We split our data into a training set (80%), a validation set (10%), and a test set (10%) in such a way that every word type was represented in all three sets of data and the number of tokens per word was proportional in all three sets. Both the LDL and the ResLDL mappings were trained on the training data and tested on the test data. In accordance with standard machine learning practice, the validation set was used to fine-tune the hyperparameters in the ResLDL model, before testing. This was not necessary for the LDL model since there are no hyperparameters in LDL. To ensure that our results were not specific to a particular data split, we repeated the entire modeling procedure 30 times, and therefore obtained thirty accuracy scores for each combination of pitch type (omnibus-segment or word F0 smooths) and network (LDL or ResLDL).
@@
We evaluated the accuracy of model predictions as follows. For each pitch vector in the test set, we obtained a corresponding predicted semantic vector and identified its closest neighbour amongst the actual CEs of the tokens in our data. If this nearest neighbour belonged to any token of the same word type as the target token, the predicted semantic vector was assessed as correct, and otherwise as incorrect. This measure of success was chosen for both computational and conceptual reasons, as detailed in the following two paragraphs.
@@
Although one might expect that a predicted semantic vector would ideally be closest in semantic space to the CE of the held-out token in question, this is computationally unrealistic. The CEs in our models are conditioned on the preceding context of a given token, and are uninformed about the following context. The pitch contours, in contrast, are shaped in part by the tone on the following word, and the probability of the word given the next word. Thus, there is information in the pitch contours that is absent in the CEs, making it computationally infeasible to predict token specific vectors. Furthermore, both the pitch contours and the CEs have measurement error. Similar to the way that a linear regression line predicts the mean value of a dependent variable for a given value of an independent variable, but not the individual data points used to generate the line, here we can predict at the level of types, but not at the level of individual tokens.
@@ 3.3.2 Results
Figure 14 presents the mean comprehension accuracies for the training data (left) and the test data (right). The individual barplots show accuracies for LDL (left two bars) and ResLDL (right two bars), for pitch contours based on segment-aware GAMs (red) and pitch contours based on word-aware GAMs (blue). As mentioned above, a given prediction is considered correct when the closest neighbour of the predicted CE is of the same word type as the target. For LDL, accuracy hovers around 30% for both training and test data, whereas for ResLDL accuracy is higher, over 50% and 40% for training and test data respectively. These results are surprisingly good, given that the models are requested to predict semantic vectors on the basis of pitch information only. For comparison, across our whole dataset the theoretical probability of a pitch vector and CE belonging to the same word type by chance is approximately 0.038. Similarly, baseline accuracies obtained by evaluating on a dataset with randomly permuted word labels were 3.7% for the training set, and 3.5% for the test set. This allows us to conclude that the classification accuracies of our models are far from trivial. On the contrary, even the least successful model achieves accuracies that are a whole order of magnitude greater than would be expected by chance.
A comparison of results from the two types of pitch vectors shows that meaning prediction is consistently more accurate when the pitch contour smooths are generated using the word GAM than using the omnibus-segment GAM. This indicates that the factor smooths for word not only contribute to a better model fit in the GAM, but also produce predicted pitch contours that are better aligned with words meanings. In other words, tonal realizations that include information about word type have the potential to help listeners to identify words’ meanings more accurately.
@@
Finally, we note that mappings from pitch contours to CEs have significant nonlinear components, as evidenced by the higher accuracies of the ResLDL model compared to the LDL model. We further note that a nonlinear mapping may be required because we are mapping from 50-dimensional pitch contours to 768-dimensional CEs. As it is mathematically impossible to map a lower-dimensional space into a higher-dimensional space with a linear mapping,202020 To see this, consider points in a cube, and their projection onto a plane in that cube. From that projection, which is two-dimensional, the original locations of the points in the three-dimensional cube cannot be reconstructed. the greater accuracy of ResLDL is unsurprising. Nevertheless, it is remarkable that the linear mappings show very similar performance on training and test data, suggesting that there is a strong linear component to predicting meaning from tonal contours.
@@
Figure 14: Mean comprehension accuracies for training data (left panel) and test data (right panel) for LDL and ResLDL mappings from omnibus-segment (red) and word (blue) pitch vectors. Mean accuracy is obtained from 30 random training and testing splits, each trained and evaluated independently. Error bars indicate double the standard error.
@@ 3.4.1 Method
Model set-up was the same as for comprehension, except that to model production the input consisted of CEs and the output consisted of pitch vectors. We again conducted the modeling procedure 30 times. For each CE in the test set, we obtained a corresponding predicted pitch vector and identified its closest neighbour amongst the actual (GAM-generated) contours of the tokens in our data. If this nearest neighbour belonged to any token of the same word type as the target token, the predicted pitch vector was assessed as correct, and otherwise as incorrect. We complemented this overall evaluation with a qualitative analysis of the pitch contours predicted by the model for individual word types. To do this, we calculated the centroid of the CEs for all tokens of a given word, and used this centroid vector to generate a predicted pitch contour from the production network with LDL mappings and word-based pitch vectors. For each of the words presented in Figure 5, we then assessed the quality of this LDL-predicted contour by visually comparing it with the contour produced by averaging the actual (GAM-generated) pitch vectors used to train the model, for all tokens of the word in question.
@@ 3.4.2 Results
Mean production accuracies (over 30 repetitions) for the token-based evaluation are presented in Figure 15. For training data (left), accuracies are between 30% and 40%. The accuracies for the test data are only slightly lower, ranging between 27% and 35%. The probability of a CE and pitch vector belonging to the same word type by chance is the same as for the comprehension models, namely 0.038. Permutation baselines are again 3.7% for training and 3.5% for testing. In other words, like the comprehension models, the production models have accuracies an order of magnitude greater than would be expected due to chance. However, in contrast to the comprehension results, production accuracies are remarkably similar for LDL and ResLDL. Apparently, linear mappings suffice when predicting low-dimensional pitch contours from high-dimensional CEs, and succeed in capturing the regularities in the meaning-to-form mappings. Possibly, predicting pitch from semantics is a cognitively more natural task than predicting semantics from just pitch on its own, and hence requires less powerful mappings. Finally, as for comprehension mappings, predicting pitch contours from CEs is more successful when pitch contours are generated with word-based GAMs, compared to segment-based GAMs.
Figure 15: Mean production accuracies for training data (left panel) and test data (right panel) for LDL and ResLDL mappings from omnibus-segment (green) and word-type (orange) pitch vectors. Mean accuracy is obtained from 30 random training and testing splits, each trained and evaluated independently. Error bars indicate double the standard error.
@@
In addition to by-word centroids, we also calculated the centroid of the CEs for all tokens of all types in our dataset. The pitch contour predicted from this overall centroid is very similar to the pitch contours in the right panel of Figure 1 and the first five panels of Figure 3. In other words, the centroid of the embeddings of all tokens can be interpreted as the ‘meaning’ of the unmodulated rise-fall pitch contour. The 10 tokens that are closest to this centroid belong to the word types {CJK*}UTF8bsmi不過 búguò ‘but, however’, {CJK*}UTF8bsmi然後 ránhòu ‘and then’, and {CJK*}UTF8bsmi時候 shíhòu ‘during which’, which suggests that these words are the most typical carriers of the RF pitch contour in the current dataset.
Figure 16: Examples of average pitch vectors (gray), and predictions generated by LDL (orange). The LDL contours were predicted from ‘centroid’ word meaning, obtained by averaging the CEs of all tokens of the same type. The word-specific contours predicted by the word GAM as presented in Figure 5, are reproduced here (after centering and scaling) in blue. The vertical dotted lines in the panels indicate the average (word-specific) syllable boundary.
@@ 4 General discussion
Our core hypothesis generates four predictions. The first prediction is that variation in tonal realization cannot be reduced to the segment-related constraints on articulation previously described in the literature. Using generalized additive models (GAMs), we were able to show that word type is indeed a more powerful predictor of tonal realization than a wide range of words’ form properties considered jointly. We not only established that GAMs with factor smooths for word type provided substantially improved model fits, but also demonstrated that word-informed GAMs provided more accurate prediction for the F0 contours of held-out data. We concluded that individual word types have specific properties — over and above their segmental makeup — that modulate the general, sine-wave shaped F0 contour characteristic of words described as having a rising-falling tone pattern. These specific properties, we conjectured, are semantic in nature.
@@
The second prediction is that information about a word’s meaning in context will improve prediction of its tonal realization. If words with the very same segments and canonical tones, but different meanings, have distinct tonal realizations, this provides evidence for the possibility that there is a semantic component to the realization of tone in Taiwan Mandarin disyllabic words. Again using generalized additive modeling, we were able to show that adding information about meaning in context (sense) led to significant improvement in model fit. On the other hand, when it came to predicting the pitch contours of unseen tokens, the ‘word’ model performed just as well as the ‘sense’ model. Having traced this equivalence to a lack of statistical power in the sense model, we ran supplementary analyses with random forests, which again revealed greater importance for the variable including sense than for word type alone. These results provide evidence for the possibility that Mandarin disyllabic word tokens indeed have tonal realizations that are partially determined by their semantics.
The third prediction is that given a pitch contour, the meaning of its carrier token can be predicted above chance level, assuming the listener has previous experience of that word type. This prediction is important for two reasons. Firstly, for meaning-specific tonal contours to facilitate speech comprehension, it is essential that pitch contours provide the listener with information that can actually be used to reduce uncertainty about a word token’s meaning. Secondly, this prediction is also motivated by a higher-level observation. Over the centuries, the phonotactics of Mandarin have been progressively simplified, resulting in a small inventory of a mere 400 different syllables. When tones are taken into account, the number of syllables increases to 2000, which is still considerably smaller than the number of syllables in use in e.g. English, which according to a perusal of the Celex database (Baayen et al., , 1995) is around 14,400. With such a restricted inventory, single-syllable Mandarin words are inevitably characterized by extensive homophony. From a functional perspective, the presence of meaning-specific pitch modulations may therefore compensate for the lack of semantic discriminability afforded by segmental makeup and syllable structure (see, e.g., Sampson, , 2015, 2019, for a discussion of theoretical implications).
@@
In order to establish empirically that token-specific tone contours indeed have the potential to help listeners zoom in on a token’s meaning, we used a specific computational modeling framework, namely the Discriminative Lexicon Model (DLM). Given the difficult task of predicting words’ high-dimensional semantic embeddings from low-dimensional pitch contours, the DLM model that we implemented performed on held-out data with an accuracy of over 50%, compared with a random baseline of 3.5%. The tonal contours of word tokens turned out to be far more revealing about their meanings than anything we thought might be possible when we started this investigation. We take these results as evidence that human listeners are presented with highly informative pitch contours, which leads us to expect that native Mandarin listeners make use of this information to optimize comprehension.
@@
The fourth and final prediction that follows from our central hypothesis is that, assuming they have previous experience of a given word type, a speaker can produce an appropriate pitch contour for a meaning they want to convey with that word. We again tested this prediction using the DLM. A network trained on matching context-specific semantic embeddings to token-specific pitch contours performed far above a random baseline, with accuracies ranging from 30% to 40% on training data, and 25% to 30% on testing data. Given that the computational models were forced to predict pitch across tokens produced by many different speakers, without any information about the segmental make-up or syllable structure of the words, this is a remarkable result that provides strong support for actual speakers in principle being able to learn to produce meaning-specific pitch contours. At this point, however, a word of caution is appropriate. Our DLM models were given the task of predicting the geometric shape of pitch contours, and did not address token- and word-specific differences in pitch height and amplitude. The modeling of the full pitch contours, including height and amplitude, is left for future investigation.
@@
Although the four predictions that follow from our central hypothesis are empirically well-supported, this of course does not necessarily imply that our hypothesis is correct. In general, given an implication from a proposition $p$ (a hypothesis) to an empirical prediction $q$, $p\Rightarrow q$, observing $q$ does not imply $p$: $q\nRightarrow p$, as there might be conditions other than $p$ that also give rise to $q$. Thus, we cannot rule out that the importance of meaning in our models might actually be due to factors that we did not take into account in our analyses. For instance, the effects of prosody, pragmatics, syntax, and emotion could, in principle, conspire to yield effects that would seem to imply token-specific semantic effects. Measures of surprisal and informativity other than the forward and backward probabilities that we included as control variables may also be informative (Tang and Shaw, , 2021). It could also be argued that in the present study, which is based on spontaneous conversational speech, the consequences of contraction and reduction (Ernestus, , 2000; Johnson, , 2004; Tseng, , 2005) are not controlled for. And indeed, we agree, all these factors are worth further investigation. However, it seems unlikely to us that any of these factors will turn out to explain away completely the effect of meaning. Our analyses are based on multiple tokens of each word type, that vary with respect to their syntactic position, their pragmatic function, the amount of segmental reduction, and their emotional valence. The pitch modulations estimated with the help of the GAM models are statistical generalizations across all this variation that is present in our data. It seems highly unlikely to us that the factors we were not able to control for will be distributed across our tokens in such an unbalanced way that they would be able to explain away our semantic effects. Furthermore, the simple fact that it is possible to predict embeddings from F0 contours, and F0 contours from embeddings, with an accuracy far beyond chance levels, bears witness to a non-trivial role for semantics in the tone system of (Taiwan) Mandarin Chinese.
It is an open question to what extent speakers and listeners actually can make use of the information that we have documented is present in token-specific pitch contours and token-specific embeddings. On the one hand, our computational model is trained under perfect conditions. This line of reasoning suggests that human learners are at a disadvantage compared to our computational model. On the other hand, human speakers and listeners encounter far more words than we considered in this study, and they understand or produce these words in much richer environments, including non-verbal aspects of communication such as gesture and facial expressions, than are available to our model. This alternative line of reasoning suggests that the human cognitive system is well-positioned for exploiting the isomorphisms between form and meaning to optimize comprehension and production.
A related question is whether listeners actually do make use of the distributional-statistical information that is in the speech signal of Mandarin RF words. We think they must be doing so, for two reasons. First, the pertinent information is present in the speech signal, but this information cannot be reduced to the consequences of bio-mechanical constraints on the speech production process. Second, the tone system of Mandarin Chinese is language specific, which argues against the possibility that the fine details of words’ pitch contours are an inevitable consequence of their meanings. Listeners must be learning the distributional statistics of tone and meaning from the speech to which they are exposed. We hasten to note that the learning of the systematicities between pitch and semantics is in all likelihood a completely subliminal process. It is not necessary for learners to be aware of the subtle modulations of pitch contours in relation to equally subtle nuances in meaning. In our conception of the learning process, successful understanding, token by token, will drive low-level learning in the lexical networks, without conscious reflection and effort being required. [^22] For token-by-token incremental lexical learning, see Heitmeier et al., 2023b , and for continuous recalibration in vision, see Marsolek, (2008).
Our central hypothesis that pitch contours in Taiwan Mandarin disyllabic words are in part determined by semantics fits well with classical studies such as those of Bybee, (2001) and Hawkins, (2003). [^23] Note that we do not claim that pitch contours are determined exclusively by semantics; the GAM analyses that we report show ample effects of non-semantic predictors. These early studies are complemented with accumulating evidence that indeed the details of phonetic realization are often intimately connected to semantics. Gahl, (2008) reported that English heterographic homophones tend to differ in duration, depending on their frequency, and Gahl and Baayen, (2024) recently showed, using distributional semantics, that differences in homophone duration can be traced in part to the meanings of these homophones. Drager, (2011) documented the ways in which the pronunciation of English ‘like’ varies with is meaning. Tomaschek et al., 2019a ’s study of the duration of English syllable-final /s/ likewise demonstrated the importance of semantics for phonetic detail.

### v2

@@ 3.2 Representing form: pitch vectors
For the DLM to implement mappings between form and meaning, every form vector input to the model has to have the same number of dimensions as all the others. However, because the tokens in our data vary in duration, our tokens also vary in the number of measurement points. This means that the raw measurements cannot be used to create the form vectors. The raw pitch contours also have the problem that there are gaps due to voicelessness. To overcome these problems, we used two of the GAMs described in Section 2 to obtain smoothed pitch contours from which we could extract a standard number of measurements. Although the sense GAM (Section 2.7) had a better fit to the data than the word GAM did (Section 2.6), the former unavoidably used a smaller dataset than the latter. For our DLM models, we wanted to maximize the number of data points available; we therefore chose to use the word model and the corresponding omnibus-segment model. We generated two predicted pitch contours for each token, one using predictions from the word GAM, and the other using predictions from the omnibus-segment GAM. Each of these predicted contours was then used to generate F0 predictions at 50 equally spaced time points ranging between 0 and 1 for every individual token.
@@
A speaker’s gender and individual characteristics such as vocal tract anatomy, idiolect, and emotional state at the time of speaking, all have strong effects both on their baseline pitch and on pitch range. Similarly, in both our word GAM and in the corresponding omnibus-segment GAM (Section 2.6), the intercepts are largely dependent on the speaker’s gender and individual identity, and differences in amplitude are largely dependent on token duration, which we take to reflect both the speakers’ idiolect and their emotional state at the time of speaking, amongst other things. On the semantic side, in normal spoken interaction between humans, a speaker’s identity and emotional state not only contribute to the pitch contours they produce, but are also conceptually available to their interlocutors. In contrast, the CEs used as semantic representations in our DLM modeling (Section 3.1) are based entirely on written text and therefore encode much less information about the speaker. To control for this discrepancy, we centered and scaled the predicted F0 values by token; that is, for each token in our data, and for each GAM, we calculated the mean and range of the 50 predicted F0 values, subtracted the mean from each predicted value, and divided the result by the range. This method of scaling (min-max normalization) ensures that the scaled data stays within a fixed range, between 0 and 1, so that every token contributes equally to the model fit, irrespective of its baseline pitch or amplitude; without scaling, tokens with a greater amplitude would be taken into account more than those with a lower amplitude. A consequence of the way we centered and scaled the pitch vectors is that our DLM production models generate predictions for the geometric shapes of the contours, but not for absolute pitch or amplitude. [^26] Geometric shape can be defined as ‘all the geometrical information that remains when location, scale, and rotational effects are filtered out from an object’ (Kendall,, 1977).
@@ 3.3.1 Method
We used two different methods to map our pitch vectors onto our semantic vectors in a comprehension network. The first method involves a straightforward linear mapping using the Linear Discriminative Learning (LDL) engine of the DLM. This is equivalent to the standard linear mappings used in statistics for multivariate multiple regression (see, e.g., Gahl and Baayen,, 2024; Heitmeier et al.,, 2021, 2025, for introductions). The second method (henceforth ResLDL) complements the linear mapping with an additional deep mapping, making it possible to accommodate nonlinear relations while keeping the model relatively interpretable. ResLDL augments an LDL mapping with a nonlinear deep network, which is given the task of capturing any systematicities that are left unexplained in the residuals of the linear network (hence the name ResLDL). Using both these methods, and comparing the results, allowed us to shed light on the complexity of the relationship between our pitch vectors and our semantic vectors. [^27] Technical details of the LDL and ResLDL implementation can be found in the online supplementary material.
@@
We split our data into a training set (80%), a validation set (10%), and a test set (10%) in such a way that every word type was represented in all three sets of data and the number of tokens per word was proportional in all three sets. In other words, the test sets consisted entirely of novel tokens but no novel types, simulating the situation for human language use with previous experience. Both the LDL and the ResLDL mappings were trained on the training data and tested on the test data. In accordance with standard machine learning practice, the validation set was used to fine-tune the hyperparameters in the ResLDL model, before testing. This was not necessary for the LDL model since there are no hyperparameters in LDL. To ensure that our results were not specific to a particular data split, we repeated the entire modeling procedure 30 times using repeated training/test splits, i.e., Monte Carlo cross-validation (Zhang,, 1993; Kuhn,, 2013). The repeated splits followed the same proportions described above, generating thirty accuracy scores for each combination of pitch type (omnibus-segment or word F0 smooths) and network (LDL or ResLDL).
@@
We evaluated the accuracy of model predictions as follows. For each pitch vector in the test set, we obtained a corresponding predicted semantic vector and identified its closest neighbor among the actual CEs of the tokens in our data. If this nearest neighbor belonged to any token of the same word type as the target token, the predicted semantic vector was assessed as correct, and otherwise as incorrect. This measure of success was chosen for both computational and conceptual reasons, as detailed in the following two paragraphs.
@@
Although one might expect that a predicted semantic vector would ideally be closest in semantic space to the CE of the held-out token in question, this is computationally unrealistic. The CEs in our models are conditioned on the preceding context of a given token, and are uninformed about the following context. The pitch contours, in contrast, are shaped in part by the tone on the following word, and the probability of the word given the next word. Thus, there is information in the pitch contours that is absent in the CEs, making it computationally infeasible to predict token-specific vectors. Furthermore, both the pitch contours and the CEs have measurement error. Similar to the way that a linear regression line predicts the mean value of a dependent variable for a given value of an independent variable, but not the individual data points used to generate the line, here we can predict at the level of types, but not at the level of individual tokens.
@@ 3.3.2 Results
Figure 13 presents the mean comprehension accuracies for the training data (left) and the test data (right). The individual barplots show accuracies for LDL (left two bars) and ResLDL (right two bars), for pitch contours based on segment-aware GAMs (red) and pitch contours based on word-aware GAMs (blue). As mentioned above, a given prediction is considered correct when the closest neighbor of the predicted CE is of the same word type as the target. For LDL, accuracy hovers around 30% for both training and test data, whereas for ResLDL accuracy is higher, over 60% and 50% for training and test data respectively. These results are surprisingly good, given that the models are requested to predict semantic vectors on the basis of pitch information only, notwithstanding the fact that pitch contains implicit information about phonological segments (cf. Section 2.6.1). For comparison, across our whole dataset the theoretical probability of a pitch vector and CE belonging to the same word type by chance is approximately 0.038. Similarly, baseline accuracies obtained by evaluating on a dataset with randomly permuted word labels were 3.7% for the training set, and 3.5% for the test set. This allows us to conclude that the classification accuracies of our models are far from trivial. On the contrary, even the least successful model achieves accuracies that are a whole order of magnitude greater than would be expected by chance.
@@
The higher accuracies of the ResLDL model compared to the LDL model indicate that mappings from pitch contours to CEs have significant nonlinear components. This nonlinear mapping may be required because we are mapping from 50-dimensional pitch contours to 768-dimensional CEs. As it is impossible to map a lower-dimensional space into a higher-dimensional space with a linear mapping without losing information,282828 To see this, consider points in a cube, and their projection onto a plane in that cube. From that projection, which is two-dimensional, the original locations of the points in the three-dimensional cube cannot be fully reconstructed. the greater accuracy of ResLDL is unsurprising. Nevertheless, it is remarkable that the linear mappings show very similar performance on training and test data, suggesting that there is a strong linear component to predicting meaning from tonal contours.
A comparison of results from the two types of pitch vectors shows that meaning prediction with ResLDL is more accurate when the pitch contour smooths are generated using the word GAM than using the omnibus-segment GAM. This indicates that the factor smooth for word not only contributes to a better model fit in the GAM, but also produces predicted pitch contours that are better aligned with words’ meanings, albeit in a nonlinear way. Of course, word encodes all the information about words’ segmental makeup that is given to the omnibus segment model, so the resulting contours do contain this information. But the superior performance of the word-based contours shows that tonal realizations that include all information associated with word type have the potential to help listeners to identify words’ meanings even more accurately than contours with just segmental information.
@@
Figure 13: Mean comprehension accuracies for training data (left panel) and test data (right panel) for LDL and ResLDL mappings from omnibus-segment (red) and word (blue) pitch vectors. Mean accuracy is obtained from 30 stratified random training and testing splits, each trained and evaluated independently. Error bars indicate double the standard error.
@@ 3.4.1 Method
Model set-up was the same as for comprehension, except that to model production the input consisted of CEs and the output consisted of pitch vectors. We again conducted the modeling procedure 30 times. For each CE in the test set, we obtained a corresponding predicted pitch vector and identified its closest neighbor among the actual (GAM-generated) contours of the tokens in our data. If this nearest neighbor belonged to any token of the same word type as the target token, the predicted pitch vector was assessed as correct, and otherwise as incorrect.
We complemented the quantitative evaluation with a qualitative analysis of the pitch contours predicted by the model for individual word types. To do this, we calculated the centroid of the CEs for all tokens of a given word, and used this centroid vector to generate a predicted pitch contour from the production network with LDL mappings and word-based pitch vectors. For each of the words presented in Figure 5, we then assessed the quality of this LDL-predicted contour by visually comparing it with the contour produced by averaging the actual (GAM-generated) pitch vectors used to train the model, for all tokens of the word in question.
@@ 3.4.2 Results
Mean production accuracies (over 30 repetitions) for the token-based evaluation are presented in Figure 14. For training data (left), accuracies are between 40% and 50%. The accuracies for the test data are only slightly lower, hovering around 40%. The probability of a CE and pitch vector belonging to the same word type by chance is the same as for the comprehension models, namely 0.038. Permutation baselines are again 3.7% for training and 3.5% for testing. In other words, like the comprehension models, the production models have accuracies an order of magnitude greater than would be expected due to chance. However, in contrast to the comprehension results, production accuracies are remarkably similar for LDL and ResLDL. Apparently, linear mappings suffice when predicting low-dimensional pitch contours from high-dimensional CEs, and succeed in capturing the regularities in the meaning-to-form mappings. Possibly, predicting pitch from semantics is a cognitively more natural task than predicting semantics from just pitch on its own, and hence requires less powerful mappings. Finally, as for comprehension mappings, predicting pitch contours from CEs is more successful when pitch contours are generated with word-based GAMs, compared to segment-based GAMs.
Figure 14: Mean production accuracies for training data (left panel) and test data (right panel) for LDL and ResLDL mappings from omnibus-segment (green) and word-type (orange) pitch vectors. Mean accuracy is obtained from 30 stratified random training and testing splits, each trained and evaluated independently. Error bars indicate double the standard error.
@@
In addition to by-word centroids, we also calculated the centroid of the CEs for all tokens of all types in our dataset. The pitch contour predicted from this overall centroid is very similar to the pitch contour in the right-hand panel of Figure 1. We therefore infer that the centroid of the embeddings of all tokens can be interpreted as the ‘meaning’ of the unmodulated rise-fall pitch contour. The 10 tokens that are closest to this centroid belong to the word types {CJK*}UTF8bsmi不過 bu2guo4 ‘but, however’, {CJK*}UTF8bsmi然後 ran2hou4 ‘and then’, and {CJK*}UTF8bsmi時候 shi2hou4 ‘during which’, which suggests that these words are the most typical carriers of the RF pitch contour in the current dataset.
Figure 15: Pitch contours for the sample of 15 word types introduced in Figure 5. The gray lines represent the average of the pitch vectors generated by the word-type GAM across all tokens of that type, i.e. the average of the contours used to train LDL. The orange lines represent the predictions generated by LDL. These LDL contours were predicted from ‘centroid’ word meaning, obtained by averaging the CEs of all tokens of the same type. The blue lines represent the word-specific contours predicted by the word GAM as presented in Figure 5 and reproduced here after centering and scaling, i.e., these blue lines show the pure effect of word on the pitch contour, irrespective of other predictors. The vertical dotted lines in the panels indicate the average (word-specific) syllable boundary.
@@ 4 General discussion
Our core hypothesis generates four predictions. The first prediction is that word type will be a stronger predictor of tonal realization than all the previously established word-form related predictors combined. This prediction follows from the hypothesis because word type includes information about meaning in addition to information about form. Using generalized additive models (GAMs), we were able to show that word type is indeed a more powerful predictor of tonal realization than a wide range of words’ form properties considered jointly. We not only established that the GAM with a factor smooth for word type provided a substantially improved model fit, but also demonstrated that the word-informed GAM provided more accurate prediction for the F0 contours of held-out data. We concluded that individual word types have specific properties — over and above their segmental makeup — that modulate the general, sine-wave shaped F0 contour characteristic of words described as having a rising-falling tone pattern. These specific properties, we conjectured, are semantic in nature.
@@
The second prediction is that information about a word’s meaning in context will improve prediction of its tonal realization. If words with the very same segments and canonical tones, but different meanings, have distinct tonal realizations, this provides evidence for the possibility that there is a semantic component to the realization of tone in Taiwan Mandarin disyllabic words. Again using generalized additive modeling, we were able to show that adding information about meaning in context, i.e., sense, did lead to significant improvement in model fit. These results provide evidence for the possibility that Mandarin disyllabic word tokens indeed have tonal realizations that are partially determined by their semantics.
@@
The third prediction is that given a pitch contour, the meaning of its carrier token can be predicted above chance level by a simple computational model with previous experience of that word type. To test this, we used the framework of the Discriminative Lexicon Model. Given the difficult task of predicting words’ high-dimensional semantic embeddings from low-dimensional pitch contours, the DLM comprehension model that we implemented performed on held-out data with an accuracy of over 50%, compared with a random baseline of 3.5%. The tonal contours of word tokens turned out to be far more revealing about their meanings than anything we thought might be possible when we started this investigation.
This finding has two important implications. Firstly, for the model to learn to predict meanings from pitch contours, pitch contours must contain information that aligns with aspects of word meaning. This suggests that human speakers of Mandarin could also potentially make use of this information to optimize speech comprehension. Secondly, meaning-specific pitch contours might be related to the extensive homophony in Mandarin. According to the Chinese Lexical Database (Sun et al.,, 2018), about 90% of monosyllabic Mandarin words have at least one homophonous counterpart. From a functional perspective, the presence of meaning-specific pitch modulations may therefore compensate for the lack of semantic discriminability afforded by segmental makeup and syllable structure (see, e.g., Sampson,, 2015, 2019, for a discussion of theoretical implications). However, given that the present study focuses on disyllabic words, which exhibit much less homophony than monosyllabic words, it remains an empirical question whether word-specific pitch contours for disyllabic words have the same functional load as contours for monosyllabic words would have. We leave this interesting question to future studies.
@@
The fourth and final prediction that follows from our central hypothesis is that, assuming it has previous experience of the relevant word type, a simple computational model can produce an appropriate pitch contour for a given meaning. We again tested this prediction using the DLM. A network trained to match context-specific semantic embeddings to token-specific pitch contours performed far above a random baseline, with accuracies ranging from 40% to 45% on training data, and 35% to 40% on testing data. Given that the computational models were forced to predict pitch across tokens produced by many different speakers, without any information about the segmental make-up or syllable structure of the words, this is a remarkable result that provides strong support for human speakers in principle being able to learn to produce meaning-specific pitch contours. At this point, however, a word of caution is appropriate. Our DLM models were given the task of predicting the geometric shape of pitch contours, and did not address token- and word-specific differences in pitch height and amplitude. The modeling of the full pitch contours, including height and amplitude, is left for future investigation.
@@
Although the four predictions that follow from our central hypothesis are empirically well-supported, this does not necessarily imply that our hypothesis is correct. We cannot rule out that the importance of meaning in our models might actually be due to factors that we did not take into account in our analyses. For instance, the effects of prosody, pragmatics, syntax, and emotion could, in principle, conspire to yield effects that would seem to imply token-specific semantic effects. Measures of surprisal and informativity other than the forward and backward probabilities that we included as control variables may also be informative (Tang and Shaw,, 2021). In addition, it could be argued that in the present study, which is based on spontaneous conversational speech, the consequences of contraction and reduction (Ernestus,, 2000; Johnson,, 2004; Tseng, 2005a, ) are not controlled for. And indeed, we agree, all these factors are worth further investigation. However, it seems unlikely to us that any of these factors will turn out to explain away completely the effect of meaning. Our analyses are based on multiple tokens of each word type, that vary with respect to their syntactic position, their pragmatic function, the amount of segmental reduction, and their emotional valence. The pitch modulations estimated with the help of the GAM models are statistical generalizations across all this variation that is present in our data. It is unlikely that the factors we were not able to control for will be distributed across our tokens in such an unbalanced way that they would be able to explain away our semantic effects. But even were it to turn out that tonal variation is completely predictable from factors other than semantics, our key point is still valid: our DLM models show that this word-specific variation is beautifully fine-tuned with words’ meanings as represented by contextualized embeddings.